A billiard ball moving at 5.0m/s strikes a stationary ball of the same mass. After the collision, the ball moves at 4.33m/s at an angle of 30.0° with respect to the original line of motion. Determine whether or not the collision was perfectly elastic.


1
Expert's answer
2022-01-12T18:08:34-0500

All Balls are identical so have same mass M.


"P_{xi}=5M\\
P_{xf}=4.33\cos{30\degree}M+ V\sin{\theta}M\\"


Applying conservation of linear momentum along original line of motion.

"V\cos{\theta}=5-4.33\times{\cfrac{\sqrt{3}}{2}}......Eq[1]""V\cos{\theta}=5-4.33\times{\cfrac{\sqrt{3}}{2}}......Eq[1]"

Applying conservation of linear momentum along perpendicular direction.

"P_{yi}=P_{yf}\\
0=4.33M\sin{30\degree}-MV\sin{\theta}\\
V\sin{\theta}= \cfrac{4.33}{2}......Eq[2]\\"

on solving Eq[1] & Eq[2], we get,

"V=\sqrt{\bigg(5-4.33\cfrac{\sqrt{3}}{2}\bigg)^2+\bigg(\cfrac{4.33}{2}\bigg)^2}\\
V=2.5m/sec""V=\sqrt{\bigg(5-4.33\cfrac{\sqrt{3}}{2}\bigg)^2+\bigg(\cfrac{4.33}{2}\bigg)^2}\\
V=2.5m/sec""V=\sqrt{\bigg(5-4.33\cfrac{\sqrt{3}}{2}\bigg)^2+\bigg(\cfrac{4.33}{2}\bigg)^2}\\
V=2.5m/sec"

and

"\theta=\tan^{-1}\bigg(\cfrac{4.33}{10-4.33\times{\sqrt{3}}}\bigg)\\
\theta=\tan^{-1}(1.72)=59.8\degree\approx60\degree""\theta=\tan^{-1}\bigg(\cfrac{4.33}{10-4.33\times{\sqrt{3}}}\bigg)\\
\theta=\tan^{-1}(1.72)=59.8\degree\approx60\degree"

yi is the same as yf so the collision was perfectly elastic since it is consistent with other data